Sipho and litha are brothers. The sum of their ages is 29. In 6 years' time Sipho will be the same age as Litha was 3 years ago. Determine the current age of the two brothers. Let x be the age of Sipho.


1
Expert's answer
2022-03-01T15:28:41-0500

Let's say sipho is X and litha is y.

Therefore "X+Y=29"

In six years time

"X+6=Y-3"

"Y=X+9"

Substitute in

"X+Y=29\\X+(X+9)=29\\sipho=10years\\ litho=19 years"